OpenStudy (anonymous):

solve: 3(4)^(3x-2)=192

jimthompson5910 (jim_thompson5910):

3(4)^(3x-2)=192 (4)^(3x-2)=192/3 4^(3x-2) = 64 4^(3x-2) = 4^3 3x - 2 = 3 ... the bases are equal (to 4), so the exponents are equal ... ... ... x = ??

OpenStudy (anonymous):

Yeah so can i marry you?

jimthompson5910 (jim_thompson5910):

lol no, sry I'm taken, but I'm flattered nonetheless I'm guessing by that response you found the value of x?

OpenStudy (anonymous):

Lol I'm taken too, I just need someone to answer math questions for me. I don't think they needed x value. Wanna answer another?

jimthompson5910 (jim_thompson5910):

yes you need to completely isolate x (ie get it by itself) ah so just marry in the math sense, lol gotcha

jimthompson5910 (jim_thompson5910):

and sure I can answer another, but tell me what you get for the value of x in this one first

OpenStudy (anonymous):

3x-2=3 3x=5 x= 5/3 Next question, \[\log _{2}(x+3)+\log_{2}(x-3)=4\]

jimthompson5910 (jim_thompson5910):

x = 5/3 is correct for the previous one

jimthompson5910 (jim_thompson5910):

\[\large \log _{2}(x+3)+\log_{2}(x-3)=4\] \[\large \log _{2}((x+3)(x-3))=4\] \[\large (x+3)(x-3)=2^4\] \[\large (x+3)(x-3)=16\] \[\large x^2 - 9 =16\] \[\large x^2 =16+9\] \[\large x^2 =25\] so x = ??? or x = ??? make sure you check your two possible answers

OpenStudy (anonymous):

okay now i have to marry you.

jimthompson5910 (jim_thompson5910):

lol I'm sure your boyfriend will be very pleased about that... jk I know you're not being serious so what two possible solutions do you get for x?

jimthompson5910 (jim_thompson5910):

and which possible solution does not work

OpenStudy (anonymous):

Positive and Negative 5, i have no clue which one doesnt work

jimthompson5910 (jim_thompson5910):

plug them in one at a time and see which one is defined and which one is not defined

OpenStudy (anonymous):

ohtay, its the negative 5. wellp. i love you.

jimthompson5910 (jim_thompson5910):

the -5 works or doesn't work?

jimthompson5910 (jim_thompson5910):

If you're not sure, let's check them one by one

jimthompson5910 (jim_thompson5910):

Let's check x = 5 \[\large \log _{2}(x+3)+\log_{2}(x-3)=4\] \[\large \log _{2}(5+3)+\log_{2}(5-3)=4\] \[\large \log _{2}(8)+\log_{2}(2)=4\] \[\large \log _{2}(8*2)=4\] \[\large \log _{2}(16)=4\] \[\large 4=4\] so that works, x = 5 is confirmed to be a true solution

jimthompson5910 (jim_thompson5910):

let's check x = -5 \[\large \log _{2}(x+3)+\log_{2}(x-3)=4\] \[\large \log _{2}(-5+3)+\log_{2}(-5-3)=4\] \[\large \log _{2}(-2)+\log_{2}(-8)=4\] this is where we run into trouble, you CANNOT take the log of a negative number...so x = -5 isn't even in the domain ---> x = -5 is NOT a solution

OpenStudy (anonymous):

I know, I meant its -5 that doesn't work lol

jimthompson5910 (jim_thompson5910):

oh gotcha